Lesson 6.1 (10 points) • Solve each proportion for x. Show your work and mark your answer clearly V+6 4
Answer #1

a oy q = -2 - 988=62-12 6 - 8 -> 6x = 9%8+12 - 6x 584 => x=84 = 14 so (x=14] va b) xt6 =4 - Qu + 56=64 9V-GV=-54 - SV=-54 ) V